OpenStudy (zyi6):

HELP ASAP!!!! Simplify. square root of 8 divided by fifth root of 8 8 to the power of negative 7 over 10 8 to the power of negative 1 over 4 8 to the power of 3 over 10 8 to the power of 7 over 10

OpenStudy (zyi6):

@hartnn

OpenStudy (zyi6):

@Nurali

OpenStudy (zyi6):

@ashhhrodriguez

OpenStudy (zyi6):



OpenStudy (zyi6):

0.2 is not a choice though, @orangesrule .

OpenStudy (nurali):

\[\frac{ \sqrt{8} }{ \sqrt[5]{8}}\]

OpenStudy (zyi6):

yes

OpenStudy (nurali):

\[\frac{ 8^{^{\frac{ 1 }{ 2 }}} }{ 8^{\frac{ 1 }{ 5 }} }\]

OpenStudy (nurali):

\[8^{\frac{ 1 }{ 2 }-\frac{ 1 }{ 5 }}\]

OpenStudy (nurali):

\[8^{\frac{ 3 }{ 10 }}\]

OpenStudy (nurali):

8 to the power of 3 over 10

OpenStudy (zyi6):

just to make sure i understand this correctly, you got the 1/2 from the invisible two above the 8 right?